OpenStudy (anonymous):

find the GCF for the following 1. 12a^5b^7c^3, 15a^2b^6c^4

OpenStudy (anonymous):

3a^2*b^6c^3

OpenStudy (anonymous):

because this expression is belonged by both the expressions...

OpenStudy (anonymous):

12a^5b^7c^3=2*2*3*a*a*a*a*a*b*b*b*b*b*b*b*c*c*c 15a^2b^6c^4=3*5*a*a*b*b*b*b*b*b*c*c*c*c

OpenStudy (anonymous):

thank you i understand these ones now